OpenStudy (anonymous):

Help Me PLEASE! 3(2x+4)-8x=-2(x+1)+3

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html this way help you step by step rule Answer: There are no solutions.

OpenStudy (anonymous):

3(2x+4)-8x = -2(x+1)+3 6x+12-8x = -2x-2+3 6x+2x-8x = -2+3 -12 -2x +2x = -14 +3 0x = -11 As BoopStars said no solution

OpenStudy (anonymous):

Perhaps you took th question down wrong?